OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches zero of quantity nine plus x divided by x to the third power. Does not exist 9 -9 0

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

plug in x=0, what do you get?

OpenStudy (anonymous):

0

OpenStudy (dumbcow):

\[\frac{9+0}{0} = \frac{9}{0} = ?\]

OpenStudy (anonymous):

so it would be does not exist?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

limit as x approaches zero of quantity negative six plus x divided by x to the fourth power.

OpenStudy (anonymous):

is that does not exist also?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

ok just checking :)

ganeshie8 (ganeshie8):

do you have options for that also ?

OpenStudy (anonymous):

i do

OpenStudy (dumbcow):

find the asymptotes lim does not exist when left limit does not equal right limit

OpenStudy (anonymous):

6,0,-6,does not exist

OpenStudy (anonymous):

can you help me with one last problem?

ganeshie8 (ganeshie8):

ok your teacher is treating "infinity" as "does not exist"

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

the other answer choices are specific, so that would make sense

ganeshie8 (ganeshie8):

yeah

OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

that was the last question

ganeshie8 (ganeshie8):

As it says, find the derivative of position function that gives you the velocity function

OpenStudy (anonymous):

ok thanks

OpenStudy (dumbcow):

wait how do you know derivatives yet if you are atill learning limits? :)

OpenStudy (anonymous):

my teacher taught them together

OpenStudy (anonymous):

that might be why I don't understand them

OpenStudy (dumbcow):

haha yeah derivatives naturally follow from limits

OpenStudy (anonymous):

well, my teacher never teaches the "normal" way.

ganeshie8 (ganeshie8):

are you able to find the derivative of position ?

OpenStudy (anonymous):

I know there is a formula for insta. velocity

ganeshie8 (ganeshie8):

interesting, what is it ?

OpenStudy (dumbcow):

hmmm...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

can you explain this to me, i cant find my binder

ganeshie8 (ganeshie8):

\[ s(t) = -2 - 6t \] Clearly s(t) is a straight line, yes ?

ganeshie8 (ganeshie8):

whats the slope of that line ?

OpenStudy (anonymous):

oh crap, is the slope -6

OpenStudy (anonymous):

I'm not thinking clearly

ganeshie8 (ganeshie8):

thats it! derivative at a point gives the slope of tangent line at that point since the position function is a straight line, the derivaitive will be same everywhere and it equals -6

ganeshie8 (ganeshie8):

thats like cheating as we did not use any calculus here, but it should be okay for now :)

OpenStudy (anonymous):

so now I have to find the velocity, right?

OpenStudy (anonymous):

or is -6 my answer?

ganeshie8 (ganeshie8):

yes the instantaneous velocity is -6 everywhere

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

thanks so much

ganeshie8 (ganeshie8):

yw